OpenStudy (anonymous):

What is the slope of the line that passes through the points (–2, 15) and (–8, –5)?

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

Plug it in the formula again. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-5-15}{-8+2}\) Simplify

OpenStudy (anonymous):

-10/3

OpenStudy (anonymous):

Um, no.. \(-5 - 15 \neq -10\) \(-8 + 2 \neq 3\)

OpenStudy (anonymous):

ur sign is wrong @zkmath

OpenStudy (anonymous):

Well, I think you simplified the fraction?

OpenStudy (anonymous):

So yeah..that's correct..but your sign is wrong..

OpenStudy (anonymous):

10/3

OpenStudy (anonymous):

Yep, you got it.